Calculate the equilibrium of income if c=R100 million / 0.8 Y and i = R125 million.
1
Expert's answer
2019-07-16T09:32:00-0400

Equlibrium level of income is attained when aggregate supply (AS) is equal to aggregate demand (AD), where AS = AD.  

Formula is Y = C + I + G

Where Y = aggregate income,C = consumption, I = investment expenditure, and G = government expenditure

For this question c=R100 million / 0.8 Y and i = R125 million

Hence equlibrium income Y= (100/0.8Y) + 125

Y=125Y + 125

Y(1-125) =125

-124Y= 125

Y= -1.0080645161